Leslie and Marnie live in two different cities that are 826 km apart. One day they both left at the same time and drove towards each other, Leslie goes 18 km/h faster than Marnie. If they met after 7 hours, how fast was each going?
Answer by mananth(16946)

let Marnie's speed be x
the distance he travels when they meet is y
y/x=7
y=7x
Leslie would have driven 826-y
speed = x+18
(826-y)/(x+18) =7
substitute y
(826-7x)/(x+18)=7
826-7x=7(x+18)
826-7x =7x+126
7x+7x=826-126
14x=700
/56
x=50
Marnie's speed = 50 km/h
Leslie's speed = 50 +!8 = 68 km/h